# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v2

@@ 5.5 Sensitivity Analysis
We study the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA. In particular, $\tau$ controls the number of SNNs propagation steps; $V_{th}^{degree}$ determines when a neuron fires; $\alpha$ governs the changing ratio of degree-aware thresholds.
Figure 4 illustrates how $\tau$, $V_{th}^{degree}$, and $\alpha$ affects the performance of DeSGDA on the PROTEINS dataset. More results on other datasets are shown in Appendix G.3. We vary $\tau$ within the range of $\{5,6,7,8,9,10\}$, $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$, and $\alpha$ in $\{0.1,0.3,0.5,0.7,0.9\}$. From the results, we observe that: (1) The performance of DeSGDA in Figure 4(a) generally exhibits an increasing trend at the beginning and then stabilizes when $\tau$ is greater than 8. We attribute this to smaller values of $\tau$ potentially losing important information for representation, while larger values significantly increase model complexity. To balance performance and complexity, we set $\tau$ to 8 as default. (2) Figure 4(b) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases. This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in the threshold, which can degrade performance. Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting the model’s ability to process information effectively. Therefore, we set $V_{th}^{degree}$ to 0.5 as default. (3) From Figure 4(c), we find that the performance of DeSGDA initially increases and then decreases as $\alpha$ increases. The potential reason is that the smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation. Contrarily, a larger $\alpha$ tends to introduce more spikes that change dynamically at each step, resulting in instability in the model’s performance. Therefore, we set $\alpha$ to 0.5 as default.
@@ 6 Conclusion
In this paper, we first propose the problem of spiking graph domain adaptation and introduce a novel framework DeSGDA for graph classification. This framework enhances the adaptability and performance of SGNs through three key aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. Our approach enables more expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively. The extensive experimental validation across benchmark datasets has demonstrated that DeSGDA not only surpasses existing methods in accuracy but also maintains efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks. In the future, we will apply SGNs in the scenarios of source-free domain adaptation and domain generalization.
Appendix A Proof of Theorem 2
@@
Proof. Before showing the designated lemma, we first introduce the following inequality to be used that:
@@
The $\hat{h}$ above is the abbreviation of $\hat{h}(x)$, which means the input is the continuous feature. Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by Yin et al. [2024], we have:
@@
Appendix B Proof of Theorem 3
@@
Proof. As proved in Theorem 2, we have:
@@ Appendix C Algorithm
Input: Source data $\mathcal{D}^{s}$; Target data $\mathcal{D}^{t}$.
Output: Node degree-aware personalized spiking graph encoder parameters $\theta$.
1:Initialize model parameters.
2:while not convergence do
3:     Sample mini-batches $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ from source and target data, respectively;
4:     Forward propagation $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through node degree-aware personalized spiking graph encoder;
5:     Pseudo-label Distilling;
6:     Calculate the loss function by Eq. 11;
7:     Update model parameters through back propagation;
@@
Appendix D Complexity Analysis
Here we analyze the computational complexity of the proposed DeSGDA. The computational complexity primarily relies on node degree-aware personalized spiking representations. For a given graph $G,\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix. $d$ is the feature dimension. $L$ denote the layer number of GIN. $|V|$ is the number of nodes. $T$ denotes the number of time latency. The spiking graph encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph. As a result, the complexity of our DeSGDA is proportional to both $|V|$ and $\|A\|_{0}$.
@@ E.1 Dataset Description
DeSGDA 68.5 72.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 70.1
Table 7: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
@@
DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.1
E.2 Data processing
@@ Appendix F Baselines
  • WL subtree: The Weisfeiler-Lehman (WL) subtree Shervashidze et al. [2011] method is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
@@
  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time..
Table 8: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance. Bold results indicate the best performance.
[…]
DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.6
DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.1
@@
  • DEAL: DEAL Yin et al. [2022] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
  • CoCo: CoCo Yin et al. [2023] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
@@
  • A2GNN: A2GNN Liu et al. [2024b] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
  • PA-BOTH: PA-BOTH Liu et al. [2024c] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
@@
Appendix G More experimental results
G.1 More performance comparison
@@
In this part, we provide additional results for our proposed method DeSGDA compared with all baseline models across various datasets, as illustrated in Table 9 to Table 17. These results consistently show that DeSGDA outperforms the baseline models in most cases, validating the superiority of our proposed method.
Additionally, the performance of DeSGDA with different GNN architectures on the NCI1, FRANKENSTEIN, and MUTAGENICITY datasets is shown in Figure 12. It is evident that GIN consistently outperforms other GNN architectures in most cases.
G.2 More Ablation study
@@
To validate the effectiveness of the different components in DeSGDA, we conduct more experiments with sive variants on NCI1, FRANKENSTEIN and MUTAGENICITY datasets, i.e., DeSGDA w/o CA, DeSGDA w/o PL, DeSGDA w/o CF, DeSGDA w/o TL, DeSGDA w PT and DeSGDA w CL. The results are shown in Table 6 , 7 and 8. From the results, we have similar observations as summarized in Section 5.4.
G.3 More Sensitivity Analysis
@@
In this part, we provide additional sensitivity analysis of the proposed DeSGDA with respect to the impact of its hyperparameters: the time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1, FRANKENSTEIN, and MUTAGENICITY datasets. The results are illustrated in Figure 9, 10 and 11, where we observe trends similar to those discussed in Section 5.5.

### v3

@@ 5.5 Sensitivity Analysis
We conduct the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA. In particular, $\tau$ controls the number of SNNs propagation steps; $V_{th}^{degree}$ determines when a neuron fires; $\alpha$ governs the changing ratio of degree-aware thresholds.
Figure 4 illustrates how $\tau$, $V_{th}^{degree}$, and $\alpha$ affects the performance of DeSGDA on the PROTEINS dataset. More results on other datasets are shown in Appendix H.4. We vary $\tau$ within the range of $\{5,6,7,8,9,10\}$, $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$, and $\alpha$ in $\{0.1,0.3,0.5,0.7,0.9\}$. From the results, we observe that: (1) The performance of DeSGDA in Figure 4(a) generally exhibits an increasing trend at the beginning and then stabilizes when $\tau$ is greater than 8. We attribute this to smaller values of $\tau$ potentially losing important information for representation, while larger values significantly increase model complexity. To balance performance and complexity, we set $\tau$ to 8 as default. (2) Figure 4(b) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases. This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in threshold, which degrades performance. Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting model’s ability to process information. Thus, we set $V_{th}^{degree}$ to 0.5 as default. (3) From Figure 4(c), we find that the performance of DeSGDA initially increases and then decreases as $\alpha$ increases. The potential reason is that smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation. Contrarily, a larger $\alpha$ tends to introduce more spikes that change dynamically at each step, resulting in instability in the model’s performance. Therefore, we set $\alpha$ to 0.5 as default.
@@ 6 Conclusion
In this paper, we propose the problem of spiking graph domain adaptation and introduce a novel framework, DeSGDA, for graph classification. This framework enhances the adaptability and performance of SGNs through three key aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. Our approach enables expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively. Extensive experimental validation across benchmark datasets demonstrates that DeSGDA surpasses existing methods in accuracy while maintaining efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks. In the future, we will apply SGNs in source-free domain adaptation and domain generalization scenarios.
Appendix A Proof of Hypothesis 1
Assuming that the node feature $h_{i}$ follows a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, then for each node in the graph, we follow the message-passing mechanism and have the information aggregation as:
$h_{i}=h_{i}+\sum_{j\in N(i)}w_{ij}h_{j}.$ (9)
Therefore, we have the expectation:
$\mathbb{E}(h_{i})=\mathbb{E}(h_{i})+\sum_{j\in N(i)}w_{ij}\mathbb{E}(h_{j}).$ (10)
Since $\mathbb{E}(h_{j})\sim\mathcal{N}(\mu,\sigma^{2})$, we have:
$\mathbb{E}(h_{i})\sim\mathcal{N}\left((1+\sum_{j\in N(i)}w_{ij})\mu,(1+\sum_{j\in N(i)}w_{ij})\sigma^{2}\right).$ (11)
From the results, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})\mu$, determined by the aggregated weights of its neighboring nodes. To provide a more intuitive understanding, we visualize the aggregated neighbor weights of GCN Kipf and Welling [2017] and GIN Xu et al. [2018] in Figure 5. The results show that as the degree increases, the aggregated weights also increase progressively. Consequently, high-degree nodes tend to follow a normal distribution with a higher mean and variance. In other words, nodes with higher degrees accumulate greater signals, making them more likely to trigger spiking. Based on this, we propose assigning higher thresholds to high-degree nodes and lower thresholds to low-degree nodes.
Another observation is that methods that normalize neighbor weights to 1 (e.g., GAT Veličković et al. [2017], GraphSAGE Hamilton et al. [2017]) still result in aggregated features following the same normal distribution. This normalization diminishes the ability to distinguish between nodes with varying degrees, ultimately degrading performance. This explains why, when using GAT as the backbone of DeSGDA, the performance is the weakest.
Figure 5: Visualization of degree-aware thresholds and aggregation weights on PROTEINS and NCI1.
Appendix B Proof of Theorem 2
@@
Proof.
Before showing the designated lemma, we first introduce the following inequality to be used that:
@@
The $\hat{h}$ above is the abbreviation of $\hat{h}(x)$, which means the input is the continuous feature. Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by Yin et al. [2024a], we have:
@@
Appendix C Proof of Theorem 3
@@
Proof.
As proved in Theorem 2, we have:
@@ Appendix D Algorithm
Input: Source data $\mathcal{D}^{s}$, Target data $\mathcal{D}^{t}$
Output: Node degree-aware personalized spiking graph encoder parameters $\theta$, domain discriminator $\gamma$
1:Initialize model parameters $\theta$ and $\gamma$
2:while not converged do
3:     Sample mini-batches $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ from $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$
4:     Forward propagate $\mathcal{B}^{s}$ and $\mathcal{B}^{t}$ through the node degree-aware personalized spiking graph encoder
5:     Perform pseudo-label distillation
6:     Compute the loss function using Eq. 8
7:     Update model parameters $\theta$ and $\gamma$ via backpropagation
@@
Appendix E Complexity Analysis
Here we analyze the computational complexity of the proposed DeSGDA. The computational complexity primarily relies on node degree-aware personalized spiking representations. For a given graph $G$, $\|A\|_{0}$ denotes the number of nonzeros in the adjacency matrix of $G$. $d$ is the feature dimension. $L$ denote the layer number of spiking graph encoder. $|V|$ is the number of nodes. $T$ denotes the number of time latency. The spiking graph encoder takes $\mathcal{O}\left(T\cdot L\cdot\left(\|A\|_{0}\cdot d+|V|\cdot d^{2}\right)\right)$ computational time for each graph. As a result, the complexity of our DeSGDA is proportional to both $|V|$ and $\|A\|_{0}$.
@@ F.1 Dataset Description
DeSGDA 68.5 71.4 70.1 69.0 68.9 66.3 69.6 70.2 71.1 69.3 74.4 70.0 69.9
Table 5: The results of ablation studies on FRANKENSTEIN (source → target). Bold results indicate the best performance.
@@
DeSGDA 63.5 64.1 62.6 63.0 81.1 62.4 65.5 65.4 81.9 60.8 82.0 65.9 68.2
Table 6: The results of ablation studies on MUTAGENICITY (source → target). Bold results indicate the best performance.
[…]
DeSGDA w CL 66.1 62.6 63.3 64.7 81.5 60.7 69.8 67.0 82.2 67.8 82.4 63.6 69.3
DeSGDA 65.4 65.9 65.5 65.6 82.8 63.6 70.7 68.2 82.9 67.6 83.9 66.5 70.8
F.2 Data processing
@@ Appendix G Baselines
  • WL subtree: Weisfeiler-Lehman (WL) subtree Shervashidze et al. [2011] is a graph kernel method, which calculates the graph similarity by a kernel function, where It encodes local neighborhood structures into subtree patterns, efficiently capturing the topology information contained in graphs.
@@
  • DRSGNN: DRSGNN Zhao et al. [2024] dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real-time.
@@
  • DEAL: DEAL Yin et al. [2022b] uses domain adversarial learning to align graph representations across different domains without labeled data, overcoming discrepancies between the source and target domains.
  • CoCo: CoCo Yin et al. [2023a] leverages contrastive learning to align graph representations between source and target domains, enhancing domain adaptation by promoting intra-domain cohesion and inter-domain separation in an unsupervised manner.
@@
  • A2GNN: A2GNN Liu et al. [2024a] introduces a novel propagation mechanism to enhance feature transferability across domains, improving the alignment of graph structures and node features in an unsupervised setting.
  • PA-BOTH: PA-BOTH Liu et al. [2024b] aligns node pairs between source and target graphs, optimizing feature correspondence at a granular level to improve the transferability of structural and feature information across domains.
@@
Table 7: Statistics of the SEED dataset.
Datasets Graphs Avg. Nodes Avg. Edges Classes
SEED 7,636 62.0 168.2 3
[…]
GCN 46.0 $\pm$ 0.9 47.8 $\pm$ 1.0 47.7 $\pm$ 1.4 49.7 $\pm$ 0.7 51.3 $\pm$ 0.8 49.8 $\pm$ 1.0
GIN 48.9 $\pm$ 0.5 50.3 $\pm$ 0.6 49.5 $\pm$ 0.7 50.7 $\pm$ 1.0 52.7 $\pm$ 1.0 52.1 $\pm$ 0.9
DEAL 53.5 $\pm$ 0.4 56.2 $\pm$ 0.7 53.2 $\pm$ 0.8 53.7 $\pm$ 1.1 55.1 $\pm$ 0.8 56.0 $\pm$ 0.7
CoCo 53.9 $\pm$ 0.5 53.0 $\pm$ 0.6 54.1 $\pm$ 0.7 54.3 $\pm$ 0.7 55.3 $\pm$ 1.0 55.9 $\pm$ 0.6
DeSGDA 54.5 $\pm$ 0.6 55.6 $\pm$ 0.7 54.6 $\pm$ 0.5 54.5 $\pm$ 0.7 55.8 $\pm$ 1.1 56.6 $\pm$ 0.8
Table 9: GPU memory consumption of different graph domain methods in training stage for each training epoch (in GB).
DeSGDA DEAL CoCo A2GNN PA-BOTH
PROTEINS 2.3 1.4 1.2 33.1 5.6
NCI1 5.7 2.9 3.3 34.8 11.9
MUTAGENICITY 5.0 2.8 3.2 35.0 12.3
FRANKENSTEIN 2.9 2.4 1.9 33.6 7.3
Table 10: Time consumption of different graph domain methods in training stage for each training epoch (in seconds).
DeSGDA DEAL CoCo A2GNN PA-BOTH
PROTEINS 0.587 0.126 22.123 0.869 0.283
NCI1 0.855 0.518 58.564 1.483 0.597
MUTAGENICITY 0.887 0.612 52.740 1.553 0.511
FRANKENSTEIN 0.663 0.375 26.837 1.016 0.275
Appendix H More experimental results
H.1 More performance comparison
@@
In this part, we provide additional results for our proposed method DeSGDA compared with all baseline models across various datasets, as illustrated in Table 1 and Table 13-23. These results consistently show that DeSGDA outperforms the baselines in most cases, validating the superiority of our proposed method. Additionally, the performance of DeSGDA with different GNN architectures on the NCI1, FRANKENSTEIN, and MUTAGENICITY datasets is shown in Figure 10. It is evident that GIN consistently outperforms other GNN architectures in most cases.
To verify the efffectiveness of DeSGDA in EEG data and multi-class scenarios, we conducted additional experiments using the SEED dataset Duan et al. [2013], Zheng and Lu [2015], which is a well-known EEG dataset for emotion classification. For EEG data processing, we utilized the torcheeg  [^5] https://github.com/torcheeg/torcheeg library to convert standard EEG data into graph structures. During graph construction, we remove edges from each graph and partitioning the dataset into source and target domains based on edge density Klepl et al. [2022]. The statistics of SEED dataset is shown in Table 7. We compared the DeSGDA with two general graph neural networks (GCN Kipf and Welling [2017] and GIN Xu et al. [2018]) and two graph domain adaptation methods (DEAL Yin et al. [2022b] and CoCo Yin et al. [2023a]). The results in Table 8 show that DeSGDA still outperforms the other methods in most cases.
H.2 Training time and memory comparison
We provide detailed comparisons of GPU memory consumption and training time per epoch for DeSGDA and other graph domain adaptation methods under identical experimental settings in this part, as shown in Tables 9 and 10. It is worth noting that the training phase is typically conducted on more powerful hardware to achieve optimal performance within a reasonable time frame.
Table 11: The results of DeSGDA with different widely used graph neural networks (GIN, GCN and SAGE) on PROTEINS and NCI1 dataset. Bold results indicate the best performance.
[…]
DeSGDA w GCN 76.6 70.5 71.8 74.1 66.3 68.3 68.5 67.1
DeSGDA w SAGE 75.8 73.3 72.4 75.2 67.2 69.5 66.6 68.4
DeSGDA w GIN 77.3 75.8 73.8 77.1 69.0 69.8 68.8 68.8
DeSGDA 78.7 78.4 74.8 77.6 68.5 71.4 70.1 69.0
Table 12: The results of DeSGDA with different widely used graph neural networks (GIN, GCN and SAGE) on MUTAGENICITY and FRANKENSTEIN dataset. Bold results indicate the best performance.
[…]
DeSGDA w GCN 60.9 60.7 63.6 60.6 61.3 62.2 60.7 61.1
DeSGDA w SAGE 61.1 62.3 64.2 61.1 61.9 62.6 61.3 61.6
DeSGDA w GIN 64.5 65.4 65.0 63.9 62.8 63.6 61.8 62.8
DeSGDA 65.4 65.9 65.5 65.6 63.5 64.1 62.6 63.0
H.3 More Ablation study
@@
To validate the effectiveness of the different components in DeSGDA, we conduct more experiments with sive variants on NCI1, FRANKENSTEIN and MUTAGENICITY datasets, i.e., DeSGDA w/o CA, DeSGDA w/o PL, DeSGDA w/o CF, DeSGDA w/o TL, DeSGDA w PT and DeSGDA w CL. The results are shown in Table 4 , 5 and 6. From the results, we have similar observations as summarized in Section 5.4.
Additionally, we conduct ablation studies to examine the effect of directly replacing the SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA: (1) DeSGDA w GCN: It replaces SGNs with GCN Kipf and Welling [2017]; (2) DeSGDA w GIN: It replaces SGNs with GIN Xu et al. [2018]; (3) DeSGDA w SAGE: It replaces SGNs with GraphSAGE Hamilton et al. [2017]. The experimental results across the PROTEINS, NCI1, MUTAGENICITY, and FRANKENSTEIN datasets are shown in Table 11 and 12. However, the critical aspect of our work lies in the specific problem we set up, i.e., low-power and distribution shift environments. In this context, directly replacing SGNs with commonly used GNNs like GIN or GCN is not feasible, as these models are unsuitable for deployment on low-energy devices. As demonstrated in Section 5.3, GNN based methods have much higher energy consumption than the spike based methods.
H.4 More Sensitivity Analysis
@@
In this part, we provide additional sensitivity analysis of the proposed DeSGDA with respect to the impact of its hyperparameters: the time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on NCI1, FRANKENSTEIN, and MUTAGENICITY datasets. The results are illustrated in Figure 11, 12 and 13, where we observe trends similar to those discussed in Section 5.5.
(a) NCI1
(b) FRANKENSTEIN
(c) MUTAGENICITY